OpenStudy (anonymous):

really need help Find the inverse of the function. f(x) = the cube root of quantity x /8. - 4

OpenStudy (anonymous):

as an example: \[ if f(x) = \sqrt[3]{x}\] Then the inverse would be \[x^3 \] So if f(x) is the cube root of (x/8-4) then the inverse is (x/8-4)^3

OpenStudy (anonymous):

f-1(x) = [8(x + 4)]3 f-1(x) = 8(x + 4)3 f-1(x) = 8(x3 + 4) f-1(x) = 24(x + 4)

OpenStudy (anonymous):

thts not in my answer choice

OpenStudy (anonymous):

I'm not sure, I think i'm missing something.

OpenStudy (anonymous):

ok can you help me withthis

OpenStudy (anonymous):

Describe how to transform the graph of f into the graph of g. f(x) = square root x-8. and g(x) = square root of quantity x+4.

OpenStudy (anonymous):

I'm not sure if this is the right form, because it's been awhile since I did this, but it's basically f(x+12) Because f(x+12) = sqrt( (x+12) - 8) = sqrt(x+4) = g(x)